# Motion Generation from Fine-grained Textual Descriptions

### v1

@@ Abstract
The task of text2motion is to generate motion sequences from given textual descriptions, where a model should explore the interactions between natural language instructions and human body movements. While most existing works are confined to coarse-grained motion descriptions (e.g., "A man squats."), fine-grained ones specifying movements of relevant body parts are barely explored. Models trained with coarse texts may not be able to learn mappings from fine-grained motion-related words to motion primitives, resulting in the failure in generating motions from unseen descriptions. In this paper, we build a large-scale language-motion dataset with fine-grained textual descriptions, FineHumanML3D, by feeding GPT-3.5-turbo with delicate prompts. Accordingly, we design a new text2motion model, FineMotionDiffuse, which makes full use of fine-grained textual information. Our experiments show that FineMotionDiffuse trained on FineHumanML3D acquires good results in quantitative evaluation. We also find this model can better generate spatially/chronologically composite motions by learning the implicit mappings from simple descriptions to the corresponding basic motions. We release our data at https://github.com/KunhangL/finemotiondiffuse.
Keywords: text2motion, fine-grained textual descriptions, prompt engineering, language-motion dataset, complex motion generation
[…]
Motion Generation from Fine-grained Textual Descriptions
Kunhang Li, Yansong Feng
Peking University
{lkh23333, fengyansong}@pku.edu.cn
Abstract content
1.   Introduction
While most works in motion generation from textual descriptions (text2motion) are devoted to designing various prediction models (Ahuja and Morency, 2019; Guo et al., 2022b; Ghosh et al., 2021; Guo et al., 2022a; Athanasiou et al., 2022; Petrovich et al., 2022; Zhang et al., 2022; Tevet et al., 2022; Kim et al., 2022; Zhang et al., 2023), the datasets they use only involve coarse-grained expressions, e.g., the short 4-word sentence in Figure 1. However, in real-world applications, we not only describe motions in high-level coarse-grained language, but also fine-grained instructions. For instance, a humanoid robot needs details about movements of relevant body parts, so as to precisely perform the target motions. However, we find that those models trained on coarse-grained descriptions cannot handle fine-grained ones. From this perspective, the available dataset itself becomes a bottleneck to improve the text2motion performance.
At the least, we expect the fine-grained descriptions to: 1) Be in time order; 2) Specify spatial changes of relevant body parts; 3) Discard unnecessary details regarding muscle tension and human feelings; 4) Conform to human body constraints. For the last expectation, a bad example is that the arms are raised in the previous sentences, but then the next sentence assumes the arms rest at both sides of the body.
@@ 1.   Introduction
To alleviate the data problem, a simple idea is to expand the coarse-grained descriptions into fine-grained ones. But it would be costly for humans to rewrite comprehensive descriptions meeting our expectations to cover a diverse range of motions. A few works attempted to utilize GPT-3 (Brown et al., 2020) to automatically generate fine-grained descriptions (Kalakonda et al., 2022; Athanasiou et al., 2023). As a large language model (LLM), GPT-3 learns much knowledge about human bodies and physical motions, which can be elicited through appropriately designed prompts. Therefore, we decide to utilize its evolutionary version, GPT-3.5-turbo  [^1] We used GPT-3.5-turbo-0301., to handle our data issue.
@@
We carefully design our prompts, test on multiple example texts and choose the best-performing one, which instructs LLM to generate fine-grained descriptions while summarizing them into pseudo-codes. We find this method greatly enhances the stability of LLM’s generation of new descriptions. With GPT-3.5-turbo, we use the selected prompt to remake HumanML3D into a high-quality language-motion dataset specializing in fine-grained texts, FineHumanML3D. To the best of our knowledge, this is the first large-scale language-motion dataset with fine-grained textual descriptions.
@@
The new fine-grained descriptions we collect are explicitly split into steps (Figure 1), but the current text2motion models do not design specific architectures to match with this property. One of the best-performing models, MotionDiffuse (Zhang et al., 2022), uses CLIP (Radford et al., 2021) with input length limit, only 77 tokens, to preprocess input texts, which appears short for most of our new descriptions. Inspired by this, we propose FineMotionDiffuse, taking fine-grained descriptions of various lengths as input, without losing any textual information due to input truncation. We also incorporate coarse-grained descriptions into FineMotionDiffuse, in order to make use of their high-level instruction-like information.
Our contributions can be summarized as follows: (1) We delicately conduct prompt engineering and evaluation, and the best prompt elicits step-based fine-grained motion descriptions, which maintain strict chronological order while specifying movements of related body parts with appropriate granularity. (2) By feeding our best-performing prompt into GPT-turbo-3.5, we build FineHumanML3D, the first large-scale language-motion dataset with fine-grained textual descriptions. (3) We propose FineMotionDiffuse, a new text2motion model simultaneously taking in our fine-grained descriptions and the original coarse-grained descriptions, to take advantage of both fine textual features encoding strict step order and coarse textual features encoding high-level short information. Human evaluation shows its strong generalization ability in dealing with complex motions.
2.   Related Work
@@
As the same motion can be described in various ways, large-scale language-motion datasets with high-quality annotations play a crucial role in text2motion. The first relevant dataset is KIT-ML (Plappert et al., 2016), but it is small in scale, with rough language descriptions and limited motion types. In comparison, HumanML3D (Guo et al., 2022b) is large in size, with a wide variety of motion types and multiple descriptions provided for each motion. However, the language descriptions of HumanML3D still tend to be coarse-grained. While there are various composite motions, their descriptions are mostly concatenations of short verb phrases that summarize each basic motion, e.g., "the figure leans down to the right, straightens, and then leans to the left".
@@
Action-GPT (Kalakonda et al., 2022) uses GPT-3 (Brown et al., 2020) to expand the original coarse-grained descriptions, but they only use rough zero-shot prompts which elicit fine-grained texts with large amounts of unnecessary details. SINC (Athanasiou et al., 2023) mines the body parts involved in multiple motion descriptions using GPT-3, and fuses their motion sequences with such information. To generate new training data, they also synthesize composite descriptions with conjunction words. They find the newly-trained model obtains the ability of spatial compositionality, that is, given a description involving multiple actions of different body parts, the model is able to generate the corresponding motion sequence simultaneously performing these actions.
@@
However, the two works do not fully explore to what extent LLM can understand, from a coarse-grained motion description, the details of time, space and human bodies. In other words, LLM’s ability to expand motion descriptions from coarse-grained to fine-grained ones still remains unknown. Moreover, we are in lack of a large-scale language-motion dataset with high-quality fine-grained descriptions. While we expect LLM to free humans from laborious work of writing new annotations, it is inevitable that we will not get perfectly aligned fine-grained descriptions and golden motions. Therefore, we are curious about whether LLM is able to complement enough descriptive details, so that the text2motion model trained on such noisy data still learns mappings from fine-grained motion-related descriptions to motion primitives.
@@
For the text2motion model, the current mainstream is to map natural language expressions and motions to the same embedding space. On the one hand, works such as TM2T (Guo et al., 2022a), TEACH (Athanasiou et al., 2022), TEMOS (Petrovich et al., 2022) and T2M-GPT (Zhang et al., 2023) use VAE-based architectures. On the other hand, benefiting from recent progress of diffusion models in generation tasks, models such as MotionDiffuse (Zhang et al., 2022), MDM (Tevet et al., 2022) and FLAME (Kim et al., 2022) successfully apply such model architectures.
@@
Of all these models, MotionDiffuse gains impressive improvements compared with other works. It utilizes the multimodal model CLIP (Radford et al., 2021), which fuses information from language and image by large-scale contrastive pretraining, to initialize its text embeddings. But CLIP only enables input texts with up to 77 tokens. Therefore, the original text encoder of MotionDiffuse cannot fully encode fine-grained descriptions. We intend to overcome this problem. Furthermore, it remains unknown whether incorporation of both coarse and fine-grained texts would help improve the text2motion model’s performance, which we aim to explore in this paper.
3.   Fine-grained Language-motion Dataset Construction
@@
We expect GPT-turbo-3.5 to properly expand motion descriptions from coarse-grained to fine-grained ones  [^2] According to our pilot study, open-source LLMs like LLaMA perform worse than GPT-turbo-3.5, so we decide to use this closed-source LLM., which depends on how we design appropriate prompts. To be more concrete, an ideal fine-grained description is supposed to state in chronological order, specify spatial changes of relevant body parts, avoid unnecessary details and conform to human body constraints. We carefully design 8 prompts and keep the best-performing one for final dataset construction. More details about these prompts and our pilot study of the prompts are in Appendix A.
3.1.   Prompt Engineering
@@
The simplest prompt, P1, is a zero-shot prompt which directly asks LLM to expand the given coarse-grained description in time order, and specify spatial position changes (including angle changes) of relevant body parts and muscle tension. The second prompt P2 does not require LLM to specify anything related to muscles, and gives one instance which includes an example coarse-grained description and its fine-grained version. P3 asks the model to give a thinking process before generating the expanded descriptions. In P4, LLM is, instead, required to explicitly mark the number of steps in order, i.e., <step n></step n>. P4 also explicitly requires LLM not to mention anything related to muscles.
The last prompt performs the best among the four. The descriptions it elicits are concise, without unnecessary details like muscle tension, finger tip movements, the agent’s feelings, etc. Benefiting from the step marks, these descriptions are in strict time order. Surprisingly, we find the involved movements highly conform to body constraints. However, this prompt is not stable enough in assigning step marks.
@@
Based on the best pilot prompt P4, we expand step marks of P5 to specify step names in angle brackets, e.g., <step 1:beginning pose>...</step 1:beginning pose>. As GPT-turbo-3.5 was trained on large amounts of codes, which is believed to increase its reasoning ability (Chen et al., 2023), we further design P6 asking LLM to summarize each step into a pseudo-code sequence. Prompt P7 and P8 are two-shot versions of P6. They share the first instance of P6, differing in that the second instance of P7 is similar to the instance of P6, while the second instance of P8 includes more code variants in the pseudo-code part.
We choose 30 coarse-grained texts to test the 4 promising prompts. Main findings are that while keeping the advantages of P4, they split steps and assign step marks more accurately. Further, we conduct a human evaluation on the quality of returned fine-grained descriptions and pseudo-codes  [^3] In all our human evaluations, evaluators are NLP researchers with adequate understanding of this task, and we reach an agreement on the evaluation standards. (Table 8 in Appendix A). Results show that the incorporation of pseudo-codes surprisingly improves the quality of returned textual descriptions. The two 2-shot prompts, P7 and P8, are proven to perform best. But a further comparison between their worse cases shows that P7 tends to elicit pseudo-codes in formats not from the instances, a sign of uncontrollability. Therefore, we select P8 as our final prompt. An example input and response elicited by P8 is shown in Figure 2.
To summarize, our key findings are: 1) LLM tends to mention muscle tension even if we do not specify this requirement, but muscles cannot be reflected in motion sequences. Therefore, we should explicitly ask LLM not to mention anything related to muscles; 2) We can require LLM to describe the motion step by step, and add named step marks to instances, which proves to be more useful than adding a general thinking process before expansion; 3) We can require LLM to further summarize the fine-grained descriptions into pseudo-codes, which increases the quality of textual descriptions; 4) We should append two diverse instances instead of one or none.
3.2.   Our FineHumanML3D Dataset
@@
The original HumanML3D dataset has 44,970 coarse-grained textual descriptions for 14,616 3D human motions. The data is augmented through motion imaging and word replacement, e.g., replacing left with right, resulting in 89,940 descriptions for 29,232 motions.
With the selected prompt, we call OpenAI’s API for GPT-3.5-turbo to expand all coarse-grained motion descriptions from HumanML3D. But GPT-3.5-turbo returns invalid responses for some of them, which we directly delete  [^4] Most are sorry-like responses, e.g., ”I’m sorry, but the description you provided is not detailed enough…”. The rest do not conform to our predefined formats, e.g., ¡p¿ marks are used instead of named step marks.. For each response, we discard the part of pseudo-codes, only keeping fine-grained descriptions with named step marks.
@@
As shown in Table 1, our FineHumanML3D is rich in vocabulary, and the average description length is nearly 10 times as that of HumanML3D. There are significantly more verbs, nouns and pronouns, indicating the appearance of more actions and body parts. Moreover, the large number of adpositions reflects more frequent interactions among body parts.
3.3.   Human Evaluation of FineHumanML3D
Although P8 shows the strong ability to elicit from GPT-3.5-turbo fine-grained descriptions with appropriate spatial details of body parts in strict chronological order, the automatically expanded texts are not guaranteed to perfectly align with the ground-truth motions. In order to assess and uncover this potential disparity, we carry out human evaluation on a sample of FineHumanML3D, i.e., 100 randomly sampled ground-truth motions along with their fine-grained descriptions.
Error / # Sub-error / #

### v2

@@ Abstract
The task of text2motion is to generate human motion sequences from given textual descriptions, where the model explores diverse mappings from natural language instructions to human body movements. While most existing works are confined to coarse-grained motion descriptions, e.g., “A man squats.", fine-grained descriptions specifying movements of relevant body parts are barely explored. Models trained with coarse-grained texts may not be able to learn mappings from fine-grained motion-related words to motion primitives, resulting in the failure to generate motions from unseen descriptions. In this paper, we build a large-scale language-motion dataset specializing in fine-grained textual descriptions, FineHumanML3D, by feeding GPT-3.5-turbo with step-by-step instructions with pseudo-code compulsory checks. Accordingly, we design a new text2motion model, FineMotionDiffuse, making full use of fine-grained textual information. Our quantitative evaluation shows that FineMotionDiffuse trained on FineHumanML3D improves FID by a large margin of 0.38, compared with competitive baselines. According to the qualitative evaluation and case study, our model outperforms MotionDiffuse in generating spatially or chronologically composite motions, by learning the implicit mappings from fine-grained descriptions to the corresponding basic motions. We release our data at https://github.com/KunhangL/finemotiondiffuse.
1 Introduction
@@
While most works in human motion generation from textual descriptions (text2motion) are devoted to designing various prediction models (Ahuja and Morency, 2019; Guo et al., 2022b; Ghosh et al., 2021; Guo et al., 2022a; Athanasiou et al., 2022; Petrovich et al., 2022; Zhang et al., 2022; Tevet et al., 2022; Kim et al., 2022; Zhang et al., 2023), the datasets they use only involve coarse-grained expressions, e.g., the short 4-word sentence in Figure 1. However, in real-world applications, we not only describe motions in high-level coarse-grained language, but also fine-grained instructions. For instance, a humanoid robot needs details about movements of relevant body parts, so as to precisely perform the target motions. We find that models trained on coarse-grained descriptions cannot handle fine-grained ones. From this perspective, available datasets exhibit a bottleneck in improving the text2motion performance.
At the least, we expect the fine-grained descriptions to: 1) Be in time order; 2) Specify spatial changes of relevant body parts; 3) Discard unnecessary details regarding muscle tension and human feelings; 4) Conform to human body constraints. For the last expectation, a bad example is that the arms are raised in the previous sentences, but the next sentence assumes that the arms rest on both sides of the body.
@@
To alleviate the data problem, a simple idea is to expand existing coarse-grained descriptions into fine-grained ones. However, it would be costly for humans to rewrite comprehensive descriptions meeting our expectations to cover a diverse range of motions. A few works attempted to utilize GPT-3 (Brown et al., 2020) to automatically generate fine-grained descriptions (Kalakonda et al., 2022; Athanasiou et al., 2023). Inspired by them, we hypothesize that large language models (LLMs) learn pertinent knowledge about human bodies and physical motions, which we aim to utilize through appropriately designed prompts  [^1] According to our pilot study, open-source LLMs like LLaMA perform worse than GPT-turbo-3.5-0301, so we decide to use this closed-source LLM..
@@
For the choice of prompts used to query the LLM for fine-grained descriptions, we manually design a set of prompt templates and evaluate their performance on multiple testing texts. The final prompt instructs the LLM to generate fine-grained descriptions while summarizing them into pseudo-codes. We find that this pseudo-code compulsory check greatly enhances the stability of the LLM’s generation of new descriptions. With this method, we utilize GPT-3.5-turbo to remake HumanML3D (Guo et al., 2022b) into a new language-motion dataset with fine-grained texts, FineHumanML3D. To the best of our knowledge, this is the first large-scale language-motion dataset specializing in fine-grained textual descriptions.
@@
In Figure 1, we exhibit an example motion with its original coarse-grained description and the newly generated fine-grained version. Our fine-grained texts are explicitly split into steps, but the current text2motion models do not design specific mechanisms to match this property. One of the best-performing models, MotionDiffuse (Zhang et al., 2022), uses CLIP (Radford et al., 2021) with an input length limit of only 77 tokens to pre-process texts, which appears short for most of our new descriptions. Inspired by this, we propose FineMotionDiffuse, which takes fine-grained descriptions of various lengths as input, without losing any textual information due to input truncation. We also incorporate coarse-grained descriptions into FineMotionDiffuse, in order to make use of their high-level instruction-like information.
Our contributions are summarized as follows:
  • By applying pseudo-code compulsory checks, our final prompt elicits step-based fine-grained motion descriptions, which maintain strict chronological order while specifying movements of related body parts with appropriate granularity.
  • By feeding our prompt into GPT-3.5-turbo, we build FineHumanML3D, the first large-scale language-motion dataset specializing in fine-grained textual descriptions.
  • We propose FineMotionDiffuse, a new text2motion model that takes in both our fine-grained descriptions and the original coarse-grained descriptions, to take advantage of both fine textual features encoding strict step order, and coarse textual features encoding high-level instruction-like information. The quantitative evaluation shows that FineMotionDiffuse trained on FineHumanML3D improves FID by a large margin of 0.38, compared with the baseline. Moreover, human evaluation shows that FineMotionDiffuse has the smallest gap from seen basic motions to unseen composite motions (0.52 compared with 0.56 for MotionDiffuse with coarse-grained descriptions and 0.80 for MotionDiffuse with fine-grained descriptions).
2 Related Work
@@
As the same motion can be described in various ways, large-scale language-motion datasets with high-quality annotations play a crucial role in solving the text2motion task. The first relevant dataset is KIT-ML (Plappert et al., 2016), but it is small in scale, with rough language descriptions and limited motion types. In comparison, HumanML3D (Guo et al., 2022b) is larger in size, with a wide variety of motion types and multiple descriptions provided for each motion. However, the language descriptions of HumanML3D still tend to be coarse-grained. While there are various composite motions, their descriptions are mostly concatenations of short verb phrases that summarize each basic motion, e.g., the figure leans down to the right, straightens, and then leans to the left.
@@
Action-GPT (Kalakonda et al., 2022) uses GPT-3 (Brown et al., 2020) to expand the original coarse-grained texts, but they only use rough zero-shot prompts that elicit fine-grained texts with large amounts of unnecessary details. SINC (Athanasiou et al., 2023) mines the body parts involved in multiple motion descriptions using GPT-3, and fuses their motion sequences with such information. To generate new training data, they further synthesize the corresponding composite descriptions with conjunction words. They find the newly-trained model obtains the ability of spatial compositionality, that is, given a description involving multiple actions of different body parts, the model is able to generate the corresponding motion sequence simultaneously performing these actions.
@@
However, the two works do not fully explore to what extent an LLM can understand, from a coarse-grained motion description, the details of time, space and human bodies. In other words, the LLM’s ability to expand motion descriptions from coarse-grained to fine-grained ones still remains unknown. Moreover, we lack a large-scale language-motion dataset with high-quality fine-grained descriptions. While we expect the LLM to free humans from the laborious work of writing new annotations, it is inevitable to get imperfectly aligned fine-grained descriptions and golden motions. Therefore, we are curious about whether the LLM is able to complement enough descriptive details, so that the text2motion model trained on such noisy data still appropriately learns mappings from fine-grained motion-related descriptions to motion primitives.
@@
For the text2motion model, the current mainstream is to map natural language expressions and motions to the same embedding space. On the one hand, works such as TM2T (Guo et al., 2022a), TEACH (Athanasiou et al., 2022), TEMOS (Petrovich et al., 2022) and T2M-GPT (Zhang et al., 2023) use VAE-based architectures. On the other hand, benefiting from the recent progress of diffusion models in generation tasks, models such as MotionDiffuse (Zhang et al., 2022), MDM (Tevet et al., 2022) and FLAME (Kim et al., 2022) successfully apply such model architectures.
@@
Of all these models, MotionDiffuse gains impressive improvements compared with previous works. It uses the multimodal model CLIP (Radford et al., 2021), which fuses information from language and images by large-scale contrastive pretraining, to initialize its text embeddings. However, CLIP only enables input texts with up to 77 tokens, so the original text encoder of MotionDiffuse cannot fully encode fine-grained descriptions. We intend to overcome this problem. Furthermore, it remains unknown whether incorporation of both coarse and fine-grained texts would help improve the text2motion model’s performance, which we aim to explore in this paper.
3 Fine-grained Language-motion Dataset Construction
@@
We expect the LLM to properly expand motion descriptions from coarse-grained to fine-grained ones. An ideal fine-grained description is supposed to state in chronological order, specify spatial changes of relevant body parts, avoid unnecessary details and conform to human body constraints. We carefully design 8 prompts and keep the best-performing one for final dataset construction. More details about these prompts and our pilot study are in Appendix A.
3.1 Prompt Engineering
@@
The simplest prompt, P1, is a zero-shot prompt that asks the LLM to expand the given coarse-grained description in time order, and specify spatial position changes (including angle changes) of relevant body parts. The second prompt P2 does not require the LLM to specify anything related to muscles, and gives one instance that includes an example coarse-grained description and its fine-grained version. P3 asks the model to give a thinking process before generating the expanded descriptions. In P4, the LLM is instead required to explicitly mark the number of steps in order, i.e., <step n></step n>. P4 further explicitly requires the LLM not to mention anything related to muscles.
The last prompt performs the best among the four. The descriptions it elicits are concise, without unnecessary details like muscle tension, fingertip movements, the agent’s feelings, etc. Benefiting from the step marks, these descriptions are in strict time order. Surprisingly, we find the involved movements highly conform to body constraints. However, this prompt is not stable enough to assign step marks.
@@
Based on the best pilot prompt P4, we expand step marks of P5 to specify step names in angle brackets, e.g., <step 1:beginning pose>...</step 1:beginning pose>. As the LLM is trained on large amounts of codes, which is believed to increase its reasoning ability (Chen et al., 2023), we further design P6 asking the LLM to summarize each step into a pseudo-code sequence. Prompt P7 and P8 are two-shot versions of P6. They share the instance of P6, differing in that the second instance of P7 is similar to the instance of P6, while the second instance of P8 includes more code variants in pseudo-codes.
We choose 30 coarse-grained texts to test the 4 promising prompts. The main findings are that while keeping the advantages of P4, they split steps and assign step marks more accurately. Further, we conduct a human evaluation on the quality of returned fine-grained descriptions and pseudo-codes  [^2] In all our human evaluations, evaluators are NLP researchers with an adequate understanding of this task, and we reach an agreement on the criteria. (Table 8 in Appendix A). Results show that the incorporation of pseudo-codes surprisingly improves the quality of returned textual descriptions. The two 2-shot prompts, P7 and P8, are proven to perform best. But a further comparison between their worse cases shows that P7 tends to elicit pseudo-codes in formats not from the instances, a sign of uncontrollability. Therefore, we select P8 as our final prompt. An example input and response elicited by P8 is shown in Figure 2.
To summarize, our key findings are: 1) The LLM tends to mention muscle tension even if we do not specify this requirement, but muscles cannot be reflected in motion sequences. Therefore, we should explicitly ask the LLM not to mention anything related to muscles; 2) We can require the LLM to describe the motion step by step, and add named step marks to instances, which proves to be more useful than adding a general thinking process before expansion; 3) We can require the LLM to further summarize the fine-grained descriptions into pseudo-codes, which increases the quality of textual descriptions; 4) We should append two diverse instances instead of one or none.
3.2 Our FineHumanML3D Dataset
@@
The original HumanML3D dataset has 44,970 coarse-grained textual descriptions for 14,616 3D human motions. The data is augmented through motion imaging and word replacement (e.g., replacing left with right), resulting in 89,940 descriptions for 29,232 motions.
With the selected prompt template, we expand all coarse-grained motion descriptions from HumanML3D to fine-grained ones. We delete invalid responses  [^3] Most are sorry-like responses, e.g., I’m sorry, but the description you provided is not detailed enough…. The rest do not conform to our predefined formats, e.g., ¡p¿ marks are used instead of named step marks.. For each response, we discard pseudo-codes and only keep fine-grained descriptions with named step marks.
@@
As shown in Table 1  [^4] We use the tokenizer and POS tagger of spaCy 3.5 (https://spacy.io/usage/v3-5) to calculate the statistics., our FineHumanML3D is rich in vocabulary, and the average description length is nearly 10 times that of HumanML3D. There are significantly more verbs, nouns and pronouns, indicating the appearance of more actions and body parts. Moreover, the large number of adpositions reflects more frequent interactions among body parts.
3.3 Human Evaluation of FineHumanML3D
Although P8 shows the strong ability to elicit from the LLM fine-grained descriptions with appropriate spatial details of body parts in strict chronological order, the automatically expanded texts are not guaranteed to perfectly align with the ground-truth motions. In order to assess and uncover this potential disparity, we carry out a human evaluation on a sample of FineHumanML3D, i.e., 100 randomly sampled ground-truth motions along with their fine-grained descriptions.
We classify all text-motion pairs into three categories — zero, partial and perfect alignment, whose counts turn out to be 2 : 68 : 30. Details of partial alignment are shown in Table 2  [^5] Each text-motion pair might have several errors.. The errors in partial alignment are categorized into inversion (e.g., left leg in the text, right leg in the motion), mismatch (partial texts not matching motion clips), redundancy (stating more than the motion performs), and deficiency (lacking descriptions of motion clips). We also show noticeable error types in redundancy and deficiency. The beginning pose type denotes the motion does not follow the beginning pose description, so for the ending pose error type. The insufficient repetition type denotes that the textual description does not repeat a single motion clip as shown by the motion, e.g., when the motion performs squatting several times, the text only describes it once.
Error / # Type / #